Find the general solution of the partial differential equation :
(D-2D’+5)(D^2+D’+3)z=e^3x+4y sin(x-y) ?
1
Expert's answer
2020-05-09T16:22:57-0400

Given differential equation is "(D-2D'+5)(D^2+D'+3)z = e^{3x}+4y sin(x-y)"

General solution z = Complement solution + Particular solution.

Auxiliary equation of given differential equation is

"(D-2D'+5)(D^2+D'+3)z = 0 \\
\implies (D-2D'+5)z = 0 \hspace{0.05 in} and \hspace{0.05 in} (D^2+D'+3)z = 0".

For "(D^2+D'+3)z = 0," assume "z = Ae^{hx+ky}" is solution,

so "Ae^{hx+ky}(h^2+k+3) =0 \implies h^2+k+3 = 0."

So complement solution(C.S.) = "e^{-5x}\phi (y+2x)+\sum Ae^{hx+ky}" where "h^2+k+3=0".


Particular solution(P.S.) of given differential equation is

P.S. = "\frac{1}{(D-2D'+5)(D^2+D'+3)} e^{3x}+4\frac{1}{(D-2D'+5)(D^2+D'+3)}y sin(x-y) \\"

"= \frac{1}{(3-2(0)+5)(3^2+0+3)} e^{3x}+ 4\frac{1}{D^3-2D^2D'+DD'+5D^2-2D'^2+3D-D'+15}ysin(x-y) \\
=\frac{1}{96} e^{3x} + A(x,y)"